OpenStudy (anonymous):

Write the standard equation of the parabola. Vertex (3,-5) and directrix x=2 Can anybody help me with this problem? I know how to solve when it is with focus instead of vertex.

OpenStudy (anonymous):

they dont tell me that

OpenStudy (anonymous):

i dont really get it;;;

OpenStudy (shamim):

is not it 3

OpenStudy (shamim):

@donghyun

OpenStudy (shamim):

3-2=1

OpenStudy (shamim):

its the distance between ur vertex n directrix

OpenStudy (shamim):

i think u r not feeling comfortable @donghyun

OpenStudy (shamim):

i think ur focus will b S(4,-5)

OpenStudy (shamim):

did u get it

OpenStudy (shamim):

distance frm focus to vertex is also 1

OpenStudy (shamim):

so the equation of parabola will b\[(x-4)^{2}+(y-(-5))^{2}=(x-2)^{2}\]